OpenStudy (anonymous):

show that any +ve odd integer is of the form 4q+1or 4q+3 where q is a +ve integer.

OpenStudy (anonymous):

If n= 2q then 2n+1 = 4q +1 If n =2q+1 2(2q+1)+1= 4q +3 Did you get it?

OpenStudy (anonymous):

Any odd integer is of the form 2n +1